OpenStudy (trojanpoem):

f(3x-4/3x+4) = x + 2 Find the equation of the tangent at the point (0, 10/3)

OpenStudy (michele_laino):

Please note that I can re-write your function as follows: \[f(x) = \frac{2}{3}\frac{{x - 5}}{{x - 2}}\]

OpenStudy (michele_laino):

oops.. I have made an error: \[f(x) = \frac{2}{3}\frac{{x - 5}}{{x - 1}}\]

OpenStudy (trojanpoem):

Well, how did you reverse it to f(x) ?

OpenStudy (michele_laino):

please you have to change variable, so please solve this equation for x: \[\frac{{3x - 4}}{{3x + 4}} = z\] where z is the new variable

OpenStudy (trojanpoem):

I see , thanks for your time.

OpenStudy (michele_laino):

thanks!